# A Survey of Text Watermarking in the Era of Large Language Models

### v3

@@ 3.3. Syntactic-based Watermarking
The lexical-based approaches attempted to embed watermark messages by replacing certain words without altering the syntax structure of the sentence. However, methods relying solely on lexical substitution are likely to lack robustness against simple watermark removal attacks, such as random synonym replacements. In the light of this, some studies attempt to embed watermarks in text in a manner that is more challenging to remove, specifically by altering the syntax structures of the text. These methods are known as syntactic-based watermarking approaches. Atallah et al. (2001) introduced three typical syntax transformations–Adjunct Movement, Clefting and Passivization–to embed watermark messages, where:
  • Adjunct Movement refers to the shifting of adjuncts to different positions within a sentence. For instance, the adverbial phrase ‘often’ can be placed in multiple locations in the sentence The dog often chased the cat.
  • Clefting is a transformational process used to emphasize a specific part of a sentence, often the subject. For example, the sentence The dog chased the cat can be transformed into It was the dog that chased the cat to highlight the dog.
  • Passivization involves converting active voice sentences with transitive verbs into the passive voice. For instance, the active sentence The dog chased the cat can be transformed into the passive form The cat was chased by the dog.
Each transformation type corresponds to a specific message bit. For instance, Adjunct Movement corresponds to 0, Clefting stands for 1, and Passivization represents 2. During watermark detection, the original and the modified text are both transformed into syntax trees. The syntax structures are subsequently compared sentence by sentence to extract watermark messages. Following this, Topkara et al. (2006a) expanded the watermark payload by additionally introducing two syntax transformation type: Activization and Topicalizaiton. Moreover, effort has not been solely made about adding watermarks into English corpus. Meral et al. (2009) investigated 20 morphosyntactic tools in Turkish. They noted that languages with high suffixation and agglutination, like Turkish, provide ample opportunities for syntactic-based watermarking.
Syntactic-based watermarking approaches can embed watermarks into existing texts in a relatively concealed manner. However, this method relies significantly on the grammatical rules of a language, potentially necessitating customization for each language. In certain texts, frequent syntactic alterations might also impact the original style and fluency of the text.
@@ 3.4. Generation-based Watermarking
Abdelnabi and Fritz (2021) introduced an end-to-end watermarking scheme named AWT. It harnesses a transformer encoder for encoding the original sentence. It then combined the sentence embedding and message embedding, feeding this composite input into a transformer decoder to derive the watermarked text. During the watermark detection process, the watermarked text is fed into transformer encoder layers to obtain the secret message. Building on top of AWT, Zhang et al. (2023a) noticed the gap between the dense distributions of the watermarked text and the sparse distributions of the one-hot watermark message encodings. To bridge this gap, they presented a watermarking method named REMARK-LLM. Likewise, its watermarking process inserts watermark message into the original text via a pretrained language model. Noticeably, a reparameterization step is introduced to transform the distribution of the generated watermarked tokens into a sparser distribution using Gumbel-Softmax (Jang et al., 2016). Then a decoder based on the transformer architecture is employed to extract the concealed messages from these embeddings. The reparameterization step allows REMARK-LLM to successfully embed two times more signatures into original text compared to prior art AWT and at the meantime maintain detection effectiveness, making a significant progress in expanding watermark payload.
@@ 4. Watermarking for LLMs
  • Step1: LLM training. The process involves training a large language model M using dataset D. The specific objectives of training can vary depending on the application context. Currently, the most prevalent training objective employed is next token prediction (Radford et al., 2019).
  • Step2: logits generation. Given a trained large language model M, a prompt $\mathbf{x}$, and a sequence of previously generated tokens $\mathbf{t}^{0:(i-1)}$, the LLM generates a probability distribution over the next token $\mathbf{t}^{(i)}$ in the vocabulary $\mathcal{V}$, represented as logits $\mathbf{l}^{(i)}$:
@@
  • Step3: token sampling. The next token $\mathbf{t}^{(i)}$ is sampled from the logits $\mathbf{l}^{(i)}$ which could be achieved using nucleus sampling (Holtzman et al., 2019), choosing the token with the highest probability (greedy decode), or using other decode algorithms such as beam search to select a list of tokens with the highest probability. Here we use $S$ to denote the token sampling process:
@@
Through these steps, the large language model M can produce a single token $\mathbf{t}^{(i)}$. To generate multiple tokens, we could simply repeat the logits generation and token sampling process iteratively.
Considering that there are three important steps during the utilization of LLM to generation text, watermarking for LLMs techniques could also be divided into three distinct types correspondingly. Specifically, we refer to the three distinct types as training time watermarking, watermarking during logits generation, and watermarking during token sampling. These three watermarking techniques will be elaborated in sections 4.1, 4.2, and 4.3, respectively.
4.1. Training Time Watermarking
The objective of training time watermarking is to embed a watermark message, denoted as $w$, into a LLM during its training phase. This process is typically accomplished by embedding the watermark message $w$ into the training dataset. Specifically, embedding a watermark message into a given dataset $D$ typically follows the following steps. The process begins with the extraction of a subset $D_{s}$ from the dataset. A watermark message is then embedded onto this subset through a watermark embedding function $W$, producing the corresponding watermarked subset $\widetilde{D}_{s}$, denoted as $\widetilde{D}_{s}=W(D_{s},w)$. Consequently, the watermarked dataset $D_{w}$ is defined as the union of the original dataset $D$ and the watermarked subset $\widetilde{D}_{s}$, minus the non-watermarked subset $D_{s}$, as expressed by the following formula:
@@ 4.1. Training Time Watermarking
Subsequently, the large language model (LLM) trained on the watermarked dataset $D_{w}$ is embedded with the watermark message $w$, denoted as $M_{w}$. Typically, $M_{w}$ exhibits characteristics of the watermark message $w$ when operating on datasets with a distribution similar to $\widetilde{D}_{s}$, enabling the wateramrk detection. In this process, the most critical aspect is the design of the watermark embedding function $W$, specifically how to transform $D_{s}$ into $\widetilde{D}_{s}$. Presently, methods for designing $W$ predominantly draw on the concept of a backdoor attack. Within the training set $\widetilde{D}_{s}=\{(x,y)\}$, for an input $x$, a specific trigger is introduced to embed a recognizable feature which manifests in the corresponding output and is detected in the subsequent verification process. Depending on whether the introduced trigger disrupts the original label $y$, the current methods can be broadly classified into two categories.
The first category of methods introduces a trigger to x, compromising the corresponding label y for that segment of x. For instance, Liu et al. (2023a) proposed a training time watermark algorithm for text classification tasks, randomly selecting a subset of text data to insert triggers at the character-level, word-level, or sentence-level. These triggers are distinct features, and they uniformly changed the labels of these data to a specific $y_{t}$. Sun et al. (2022) implemented a similar approach for code generation, applying word-level or sentence-level triggers to text, and employing methods like code corrupting to disrupt the associated code. Although these techniques are effective for detecting models using trigger-inclusive text, they compromise part of the dataset’s labels, potentially degrading the model’s inherent capabilities.
To address this issue, methods that produce low distortion on labels are required. Tang et al. (2023) initially adopt adversarial learning to identify examples within a certain category C that are prone to be wrongly classified by models and then add triggers to these samples without altering their original labels. For code generation tasks, Sun et al. (2023) conduct semantically invariant transformations of the code, such as employing different forms of syntactic sugar. This allows for the matching and detection of triggers in text with varying code styles. Currently, the objective of training time watermarking is to protect the copyright of datasets from unauthorized use.
Although Training Time Watermarking can embed watermark messages into large language models, it has several distinct disadvantages. Firstly, the watermarked LLM $M_{w}$, can only produce watermarked outputs for a subset of inputs. Secondly, the watermark payload is considerably limited, typically only indicating the presence of a watermark and not containing more extensive information. Thirdly, the watermark message is challenging to modify: altering the watermark message $w$ often requires retraining the model, which is a costly process. Therefore, the application of Training Time Watermarking remains quite limited.
@@ 4.2. Watermarking during Logits Generation
Watermarking during logits generation refers to the insertion of a watermark message $w$ into the logits (i.e., probabilities of tokens in the vocabulary $\mathcal{V}$) generated by large language models. This approach does not necessitate alterations to the parameters of the LLM $M$, rendering it more flexible and cost-effective compared to training time watermarking methods.
In the scenario of watermarking during logits generation, the watermarking algorithm $\mathcal{A}$ modifies the logits produced by the LLM. Specifically, the watermark’s message $w$ is embedded into the logits generated by LLM. The watermarked logits $\widetilde{\mathbf{l}^{(i)}}$, could be calculated using the following formula:
@@
where we assume the watermarked logits $\widetilde{\mathbf{l}^{(i)}}$ is generated by a watermarked LLM $M_{w}$.
Kirchenbauer et al. (2023a) proposed the first watermarking method based on LLM logits modification, referred to as KGW. There, the watermark generator randomly splits the vocabulary set into a red list and a green list at each position depending on the previous token using a hash function. As illustrated in Figure 4, when $M_{w}$ generates the $i^{th}$ token, a small bias $\delta$ is added to the logits of the tokens in the green list. Let $G$ represent the green list and $R$ stands for the red list. Then the logits value of token $v_{j}$ at position $i$ would be calculated as follows:
@@
As the watermark algorithm favors the logits of green tokens, the watermarked text is expected to contain a higher proportion of green tokens compared to the unwatermarked text. Consequently, the watermark detector assesses whether a text is watermarked by first utilizing the hash function to categorize each token as red or green sequentially. Following this, it calculates the green token proportion using the z-metric. If the green token proportion surpasses a specific threshold, the text is classified as watermarked.
@@
Although the KGW watermark detection method demonstrated exceptional performance in its test scenarios, achieving a false positive rate (human text misclassified as watermark text) of less than $3\times 10^{-3}$%, and a false negative rate (watermark text not detected) of less than 1%, there remain numerous highly challenging scenarios in real-world applications that necessitate specialized optimization and design of the watermark algorithm to effectively address them. Below, we list four representative scenarios and thoroughly introduce the improvements and explorations made to watermark algorithms in these contexts, which are illustrated in Figure 5.
@@
Low-entropy scenarios are situations such as code generation and formatted document generation, where the generated text is relatively deterministic. Entropy serves as a quantification of textual uncertainty, which is calculated as follows:
@@
where $P_{j}^{(i)}$ denotes the probability value of token $v_{j}$ at position $i$. A decrease in entropy corresponds to a heightened level of certainty within the produced text. In real applications, it is also necessary to add watermark when using LLM to generate low-entropy text. Under low-entropy scenarios, significant modifications are not allowed, since it may notably degrade text quality. Therefore, watermarking and detecting watermark in low-entropy text pose a challenging task.
To address this problem, Lee et al. (2023) suggested calculating the entropy before adding preference to logits of green tokens (i.e. set G in Equation 9). If the entropy $H^{(i)}$ is lower than a threshold $H$, the logits vector remains unmodified. Similarly, Wang et al. (2023b) employed a balance-marking strategy, constraining the choice of vocabulary to a subset where the cumulative model log probabilities exceed a certain percentage. This approach effectively allows for the watermarking on only high-entropy tokens while bypassing low-entropy tokens. For instance, consider a specific case where there is only one word candidate; in such a scenario, the vocabulary subset is limited to this single candidate. The above methods (Lee et al., 2023; Wang et al., 2023b), by considering the influence of text entropy during the watermarking process, achieves a relatively low impact on the quality of the text. However, this method is still ineffective when the number of high-entropy tokens in the text is low. Adding watermarks to low-entropy text remains a challenge.
@@
The KGW(Kirchenbauer et al., 2023a) watermark algorithm introduced by Equation 9, could only determine the presence of a watermark and no additional information, categorizing it as a zero-bit payload watermark. However, practical applications often demand that watermarks carry additional information such as copyright details, timestamps, or specific identifiers. This necessitates watermarking techniques that not only detect the presence of watermarks but also extract meaningful information, which could be categorized as multi-bit payload watermark.
To achieve this, one possible solution is to establish multiple splits for dividing the vocabulary into a red list and a green list. Specifically, we can set N types of splits for the vocabulary, such as $(G_{1},R_{1}),…,(G_{N},R_{N})]$, where each split could performing LLM watermarking following the way introduced in Equation 9. Each split corresponds to a specific watermark message, endowing the watermark with a $log_{2}{N}$-bit payload. For instance, Wang et al. (2023b) permits the user to input a watermark message $m$ with $log_{2}{N}$ bits. Subsequently, based on the hash value of this message $m$, the vocabulary is divided. However, during detection phase, this method requires iterating through all $N$ possible messages, calculating the correlation between the text and the red/green splits corresponding to each message respectively, thus is not efficient. To mitigate this issue, Fernandez et al. (2023) proposed the Cyclic Shift method, where distinct messages are generated by cyclically shifting an initial message. This approach reduces redundant computational steps and enhances detection efficiency through parallel processing.
However, these methods, due to the necessity of iterating each possible message, encounter heightened computational complexity with the growing watermark payload. To address this, Yoo et al. (2023a) proposed allocating different bits of the message to various positions in the watermark text, allowing independent detection of different bit altogether in a single detection round. Then the information corresponding to each bit is concatenated to obtain the final message. Specifically, the message is modeled as a sequence $m=\Sigma^{b}$ where $\Sigma=\{0,1\}$. The process of dividing the vocabulary at each generation step is then split into two phrases: the first involves choosing a position in the message, $\Sigma^{b}[i]$, based on the existing random seed, and the second involves dividing the vocabulary according to this $\Sigma^{b}[i]$ and random seed. This approach significantly improves efficiency by enabling the parallel detection of bit information. Furthermore, Yoo et al. (2023a) also suggest dividing the vocabulary into more parts, so that each position $\Sigma^{b}[i]$ in $m$ can contain more information, further enhancing the payload of the text watermark.
@@
As discussed in section 2, an effective watermarking algorithm must possess sufficient robustness against watermark removal attacks, ensuring that the watermark text remains detectable post-attack. These attacks typically involve modifications to the text without altering its semantic content, which will be introduced in section 5.3. Although the KGW algorithm (Kirchenbauer et al., 2023a) demonstrated some robustness to watermark removal attacks in their experiments, there remains room for improvement in its robustness.
In particular, for the watermarking during logits generation methods (Kirchenbauer et al., 2023a, b; Zhao et al., 2023a; Liu et al., 2023b, c), the most significant factor affecting robustness is how to determine the modifications to logits. More specifically, how to divide the red-green list mentioned in the KGW method (Kirchenbauer et al., 2023a). In the original KGW method (Kirchenbauer et al., 2023a), the red-green list is determined based on the hash value of preceding tokens. Kirchenbauer et al. (2023b) further elaborated on some specific hash strategies, such as using only the token with the smallest token id in previous tokens for hashing to decide the red-green list, which can enhance robustness. (Zhao et al., 2023a) proved that a globally fixed split of red and green list results in higher robustness to watermark removal attacks . Additionally, since watermark removal attacks typically do not alter the semantic content of the text, many studies have also designed methods to determine the split of the red-green list based on textual semantics. For example, Liu et al. (2023c) trained a watermark model that can directly convert text semantic embeddings into watermark logits. Ren et al. (2023) converted semantic embeddings into semantic values through weighted embedding pooling followed by discretizing using NE-Ring, and then divided the vocabulary into red-list and green-list based on these semantic values. The current methods primarily focus on investigating the robustness of zero-bit watermark algorithms against watermark removal attacks. Future algorithms could further explore the robustness of multi-bit payload watermark algorithms.
@@
In the aforementioned discussion on the watermark algorithms’ ability to preventing watermark removal attacks, it is assumed that the attacker is not aware of the details and generation methods of the watermark algorithm, which is essentially a black-box setting. Once an attacker obtain the watermark generation details, they could effortlessly remove the watermark using the anti-watermark techniques (Kirchenbauer et al., 2023b) or forge the watermark easily. Therefore, for watermark algorithms, the capability to defend against watermark forgeries is of great importance. The ability to defend against watermark forgeries depends on the watermark algorithm’s capacity to effectively conceal its watermark generation process.
If a watermark algorithm produces watermarked text that is imperceptible, the watermark would become more difficult to forge. Here, imperceptible refers to the indistinguishability in distribution between watermarked and non-watermarked texts. Hu et al. (2023) discovered that the way KGW algorithm (Kirchenbauer et al., 2023a) modifies to logits are biased, thus lacking the imperceptibility. Specifically, biased in this context is defined as the expectation of watermarked logits for all keys being the original logits of the language model:
@@
where each distinct key corresponds to a different method of splitting the red-green list. The key reason why the KGW algorithm is biased is that it applies a uniform bias, $\delta$, to all tokens in the green list. However, this uniform $\delta$ disproportionately impacts tokens with lower probabilities, ultimately resulting in bias (as detailed in the proof by (Hu et al., 2023)). To address this issue, Hu et al. (2023) proposed two reweighting methods to make the watermarking algorithm unbiased. The $\delta$-reweight method samples a one-hot distribution directly based on the original logits’ distribution. In contrast, the $\gamma$-reweight method randomly rejects half of the probability distribution range, doubling the probabilities of the remaining tokens. Similarly, Wu et al. (2023) employed an $\alpha$-reweight method, which rejects tokens with probabilities below $\alpha$ and proportionally increases the probabilities of the rest. Theoretically, these three algorithms can be proven to be unbiased. These unbiased watermarking algorithms, compared to the KGW algorithm, are better at being imperceptible. However, the unbiased distribution resulting from these re-weighting methods may not guarantee theoretical imperceptibility. For instance, the variance in the distributions with same mathematical expectations might be distinct. Therefore, further work is required to explore whether these algorithms can truly achieve imperceptibility.
@@
The capability of watermark algorithms to defend against watermark forgeries cannot be solely based on the imperceptibility of the watermark. It is also crucial that these algorithms robustly protect watermark rules from being deciphered. In this context, we differentiate between two scenarios: private detection and public detection. Private detection refers to cases where watermark detection is only accessible to users through an API. In contrast, public detection implies that the detail of watermark detector is open to all users. For private detection scenarios, the complexity of the watermark algorithm plays a vital role in its ability to resist watermark forgeries. For instance, Zhao et al. (2023a) adopted a fixed global red-green list split to generate watermark logits. Such a simple rule can be easily cracked through statistical analysis of the watermark text (Sadasivan et al., 2023), revealing which tokens are included in the green token list. Conversely, Liu et al. (2023c) employed the semantic information to generate watermarked logits. Extracting the watermark rules from their produced watermark text is considerably more challenging, as these rules vary with different texts and are sufficiently complex.
@@
In the context of public detection scenarios, resisting watermark forgeries becomes significantly more challenging. This difficulty arises because attackers can access the watermark detectors. For methods where the watermark generation process is involved in detection (Kirchenbauer et al., 2023a, b; Zhao et al., 2023a; Liu et al., 2023c), exposing the watermark generator leads to a complete inability to resist watermark forgeries. To address this issue, Fairoze et al. (2023) have utilized digital signature technology from the field of cryptography. This approach involves generating watermarks using a private key and verifying them with a public key. However, the verification via public key relies on features extracted from the text and users can still exploit these features to forge watermarks. Further advancing this field, Liu et al. (2023b) proposed the use of neural networks for watermark detection. Due to the black-box nature of neural networks, the details of watermark generation are not exposed, which could defend against watermark forgeries in public detection scenarios.

### v4

@@ 3.3. Syntactic-based Watermarking
The lexical-based methods aim to embed watermarks by substituting specific words, maintaining the sentence’s syntax. Yet, these approaches, relying exclusively on lexical substitution, might not be robust against straightforward watermark removal tactics like random synonym replacement. Consequently, several studies have explored embedding watermarks in a manner that resists removal, notably by modifying the text’s syntax structure. These methods are known as syntactic-based watermarking approaches. Atallah et al. (2001) introduced three typical syntax transformations–Adjunct Movement, Clefting and Passivization–to embed watermark messages, where:
  • Adjunct Movement entails relocating adjuncts within a sentence. Example: ‘often’ can be variably positioned in The dog often chased the cat.
  • Clefting emphasizes a sentence part, typically the subject. For instance, The dog chased the cat becomes It was the dog that chased the cat to accentuate the dog.
  • Passivization changes active sentences with transitive verbs to passive voice. For example, The dog chased the cat is passivized to The cat was chased by the dog.
Each transformation type is assigned a unique message bit: Adjunct Movement to 0, Clefting to 1, and Passivization to 2. In watermark detection, both original and altered texts are converted into syntax trees, and their structures are compared for message extraction. Expanding this concept, Topkara et al. (2006a) introduced additional syntax transformations: Activization and Topicalization. Moreover, research extends beyond English, with Meral et al. (2009) analyzing 20 morphosyntactic tools in Turkish, highlighting that languages with significant suffixation and agglutination, such as Turkish, are well-suited for syntactic watermarking.
While syntactic-based watermarking effectively embeds watermarks in a concealed manner, it heavily depends on a language’s grammatical rules, often requiring language-specific customization. Frequent syntactic changes in some texts may also alter their original style and fluency.
@@ 3.4. Generation-based Watermarking
Abdelnabi and Fritz (2021) developed AWT, an end-to-end watermarking scheme utilizing a transformer encoder to encode sentences and merge sentence and message embeddings. This composite is processed by a transformer decoder to generate watermarked text. For watermark detection, the text undergoes transformer encoder layers to retrieve the secret message. Extending AWT, Zhang et al. (2023a) identified disparities between the dense watermarked text distributions and sparse one-hot watermark encodings. Addressing this, they proposed REMARK-LLM, employing a pretrained language model for watermark insertion. A key innovation is the reparameterization step using Gumbel-Softmax (Jang et al., 2016) to yield sparser token distributions. A transformer-based decoder extracts messages from these embeddings. REMARK-LLM notably embeds double the signatures of AWT while maintaining detection efficacy, significantly enhancing watermark payload capacity.
@@ 4. Watermarking for LLMs
  • Step1: LLM Training. This step involves training a large language model, M, with a dataset D. Training objectives vary based on the application, with next token prediction being the most common (Radford et al., 2019).
  • Step2: Logits Generation. With the trained LLM M, given a prompt $\mathbf{x}$ and a sequence of prior tokens $\mathbf{t}^{0:(i-1)}$, the LLM predicts the next token $\mathbf{t}^{(i)}$’s probability distribution in the vocabulary $\mathcal{V}$, expressed as logits $\mathbf{l}^{(i)}$:
@@
  • Step3: Token Sampling. The next token $\mathbf{t}^{(i)}$ is selected from $\mathbf{l}^{(i)}$, using methods like nucleus sampling (Holtzman et al., 2019), greedy decoding, or beam search. The sampling process is denoted as:
@@
Through these steps, LLM M generates a token $\mathbf{t}^{(i)}$. For multiple tokens, logits generation and token sampling are iteratively repeated.
In aligning with the three critical phases of text generation using Large Language Models (LLMs), watermarking techniques for LLMs are similarly categorized into three distinct types. These are: watermarking at training time, during logits generation, and in the token sampling phase. Detailed discussions of these watermarking methods are presented in sections 4.1, 4.2, and 4.3, respectively.
4.1. Watermarking during LLM Training
The objective of training time watermarking is to insert a watermark message, $w$, into a Large Language Model (LLM) during training phase. This involves incorporating $w$ into the training dataset $D$. The procedure starts with selecting a subset $D_{s}$ from $D$. The watermark message is embedded in $D_{s}$ using a watermark embedding function $W$, resulting in the watermarked subset $\widetilde{D}_{s}$, where $\widetilde{D}_{s}=W(D_{s},w)$. The watermarked dataset $D_{w}$ is then constructed by combining the original dataset $D$ with $\widetilde{D}_{s}$, after removing $D_{s}$:
@@ 4.1. Watermarking during LLM Training
An LLM trained on $D_{w}$ integrates the watermark $w$, yielding a model $M_{w}$. This model displays characteristics of $w$ when used on data resembling $\widetilde{D}{s}$, facilitating watermark detection.
The critical element in this process is designing the embedding function $W$, particularly transforming $D_{s}$ into $\widetilde{D}_{s}$. Current strategies for crafting $W$ are largely inspired by backdoor attack concepts, where a specific trigger in the training set $\widetilde{D}{s}={(x,y)}$ is embedded in input $x$ to create a detectable output feature. Depending on the trigger’s impact on the original label $y$, these methods are broadly categorized into two types.
The first category involves inserting a trigger into x, altering the corresponding label y. For example, Liu et al. (2023a) introduced a watermark algorithm for text classification, inserting triggers at character, word, or sentence levels in a text subset, uniformly changing their labels to a specific $y_{t}$. Sun et al. (2022) applied a similar strategy in code generation, using word or sentence-level triggers and code corrupting. While these methods effectively detect models with trigger-incorporated text, they may degrade the model’s performance by compromising label integrity.
To mitigate label distortion, alternative approaches are employed. Tang et al. (2023) used adversarial learning to identify and trigger-mark samples within category C without changing their labels. For code generation, Sun et al. (2023) performed semantically invariant code transformations, such as syntactic sugar variations, facilitating trigger detection in varied code styles. The primary goal of training time watermarking is to safeguard dataset copyrights against unauthorized use.
Despite its ability to embed watermarks in LLMs, Training Time Watermarking has notable drawbacks. The watermarked LLM $M_{w}$ generates watermarked outputs only for specific inputs. The watermark payload is limited, mainly serving as an indicator rather than conveying extensive information. Modifying the watermark message $w$ usually necessitates model retraining, a resource-intensive task. Hence, the application scope of Training Time Watermarking is relatively restricted.
@@ 4.2. Watermarking during Logits Generation
Watermarking during logits generation refers to the insertion of a watermark message $w$ into the logits generated by LLMs. This technique, which does not require modifying the LLM parameters, is more versatile and cost-effective than training time watermarking methods.
In this context, the watermarking algorithm $\mathcal{A}$ alters the logits from the LLM to incorporate the watermark message $w$. The modified logits $\widetilde{\mathbf{l}^{(i)}}$ can be computed as follows:
@@
where $\widetilde{\mathbf{l}^{(i)}}$ is assumed to be produced by a watermarked LLM $M_{w}$.
Kirchenbauer et al. (2023a) introduced the first LLM watermarking technique based on logits modification, termed KGW. This method partitions the vocabulary into a red list (R) and a green list (G) at each token position, using a hash function that depends on the preceding token. As depicted in Figure 4, for the $i^{th}$ token generation by $M_{w}$, a bias $\delta$ is applied to the logits of tokens in G. The adjusted logit value, $\widetilde{\mathbf{l}_{j}^{(i)}}$, for a token $v_{j}$ at position $i$ is calculated as follows:
@@
This algorithm biases towards green tokens, leading to a higher proportion in watermarked texts. The detector categorizes each token as red or green using the hash function and calculates the green token ratio with the z-metric, defined as:
(10) $z=(\lvert s\rvert_{G}-\gamma T)/\sqrt{T\gamma(1-\gamma)}$
where T is the length of the text, $\gamma$ is the ratio of the green list. A text exceeding a certain green token threshold is deemed watermarked.
@@
KGW’s detection method showed low false positive (< $3\times 10^{-3}$%) and false negative (< 1%) rates in tests. Yet, real-world application challenges necessitate further optimization and design. The following outlines four representative scenarios, with improvements and explorations in watermark algorithms, depicted in Figure 5.
@@
In low-entropy scenarios, such as code or formatted document generation, text is typically deterministic. Entropy, measuring textual uncertainty, is calculated as follows, where $P_{j}^{(i)}$ denotes the probability of token $v_{j}$ at position $i$:
@@
Lower entropy implies higher text predictability. Watermarking in such contexts is challenging due to the need of minimal text alterations.
Lee et al. (2023) proposed calculating entropy before modifying logits of green tokens (set G in Equation 9). If $H^{(i)}$ falls below a threshold $H$, logits are kept unchanged. Likewise, Wang et al. (2023b) introduced a balance-marking strategy, limiting vocabulary choices to subsets with cumulative log probabilities above a specific level. This method focuses on watermarking high-entropy tokens while avoiding low-entropy ones. For example, in cases with a single word candidate, the subset is restricted to this candidate. These methods (Lee et al., 2023; Wang et al., 2023b), by factoring in text entropy, minimize impact on text quality. However, watermarking remains problematic in texts with few high-entropy tokens.
@@
The KGW watermark algorithm (Kirchenbauer et al., 2023a), as defined in Equation 9, can only verify watermark presence, classifying it as a zero-bit watermark. Yet, many applications require watermarks to convey additional information like copyright details, timestamps, or identifiers, leading to the need for multi-bit watermarks capable of extracting meaningful data.
To achieve this, one possible solution involves multiple vocabulary divisions into red and green lists, creating N types of splits $(G_{1},R_{1}),…,(G_{N},R_{N})]$. Each split corresponds to a distinct watermark message, offering a $log_{2}{N}$-bit payload. For example, Wang et al. (2023b) allows inputting a $log_{2}{N}$-bit watermark message $m$, with the vocabulary division based on the hash value of $m$. However, this approach is inefficient during detection, requiring iteration over all $N$ messages to calculate text correlation with each split. To address this, Fernandez et al. (2023) proposed the Cyclic Shift method to enhance efficiency, generating different messages by cyclically shifting an initial one, allowing parallel processing.
Nonetheless, these methods face increased computational complexity with larger watermark payloads due to the necessity of iterating each possible message. To mitigate this, Yoo et al. (2023a) suggested assigning different message bits to distinct text positions, facilitating simultaneous bit detection. The message $m$ is modeled as a sequence $m=\Sigma^{b}$ with $\Sigma=\{0,1\}$. Vocabulary division at each step involves selecting a message position $\Sigma^{b}[i]$ based on the existing random seed, then dividing the vocabulary based on the selected $\Sigma^{b}[i]$. This method improves efficiency by enabling parallel bit detection. Additionally, Yoo et al. (2023a) recommended dividing the vocabulary into more segments, so each position $\Sigma^{b}[i]$ in $m$ holds more data, increasing the watermark’s payload.
@@
As discussed in section 2, an effective watermarking algorithm must possess sufficient robustness against watermark removal attacks, ensuring that the watermark text remains detectable after an attack. These attacks typically involve modifications to the text without altering its semantic content, which will be introduced in section 5.3. Although the KGW algorithm (Kirchenbauer et al., 2023a) demonstrated some robustness to watermark removal attacks in their experiments, there remains room for improvement in its robustness.
For watermarking during logits generation (Kirchenbauer et al., 2023a, b; Zhao et al., 2023a; Liu et al., 2023b, c), the pivotal factor for robustness lies in the methodology used for modifying logits, especially the red-green list division in KGW (Kirchenbauer et al., 2023a). The original KGW utilizes the hash values of preceding tokens for this division. Kirchenbauer et al. (2023b) further elaborated on some specific hash strategies, such as using only the token with the smallest token id in previous tokens for hashing to decide the red-green list, which can enhance robustness. Zhao et al. (2023a) proved that a fixed global split of red and green lists offers greater resistance to removal attacks. Additionally, as watermark removal attacks usually preserve the semantic content of the text, several studies have developed methods to determine the red-green list split based on textual semantics. For example, Liu et al. (2023c) trained a watermark model that can directly convert text semantic embeddings into watermark logits. Ren et al. (2023) converted semantic embeddings into semantic values through weighted embedding pooling followed by discretizing using NE-Ring, and then divided the vocabulary into red-list and green-list based on these semantic values. The current methods primarily focus on investigating the robustness of zero-bit watermark algorithms against watermark removal attacks. Future algorithms could delve into exploring the robustness of multi-bit payload watermark algorithms.
@@
In the previously discussed watermark algorithms’ resilience to removal attacks, it is assumed that the attacker operates in a black-box scenario, lacking knowledge about the details and generation methods of the watermark. If an attacker acquires this information, removing or forging the watermark becomes trivial (Kirchenbauer et al., 2023b). Consequently, the ability of watermark algorithms to withstand forgeries is crucial. The ability to defend against forgeries depends on the watermark algorithm’s capacity to effectively conceal its watermark generation process.
If a watermark algorithm yields imperceptible watermarked text, where imperceptible means indistinguishable distribution between watermarked and unwatermarked texts, it becomes more challenging to forge. Hu et al. (2023) noted that the KGW algorithm (Kirchenbauer et al., 2023a) introduces bias in its logits modification, compromising imperceptibility. Imperceptibility here implies that the expected watermarked logits equal the original model’s logits:
@@
where each key $k$ represents a unique red-green list split. The bias in KGW arises from applying a uniform $\delta$ to green list tokens, disproportionately affecting low-probability tokens, leading to bias. To counter this, Hu et al. (2023) introduced two unbiased reweighting methods: $\delta$-reweight, sampling a one-hot distribution from original logits, and $\gamma$-reweight, which halves the probability distribution range, doubling the remaining tokens’ probabilities. Similarly, Wu et al. (2023) proposed the $\alpha$-reweight method, discarding tokens with probabilities below $\alpha$ and adjusting the rest. These methods, theoretically unbiased, improve imperceptibility over KGW. However, unbiased distributions may not ensure imperceptibility, as variances in distributions with identical expectations can differ. Further research is needed to ascertain true imperceptibility in these algorithms.
@@
The capability of watermark algorithms to defend against watermark forgeries cannot be solely based on the imperceptibility of the watermark. It is also crucial that these algorithms robustly protect watermark rules from being deciphered. In this context, we differentiate between two scenarios: private detection and public detection. Private detection allows watermark verification solely via an API, whereas public detection involves open-access watermark detector details. In private scenarios, the watermark algorithm’s complexity is critical for preventing forgeries. For example, Zhao et al. (2023a) used a fixed red-green list split for watermark logits, a simplistic approach vulnerable to statistical analysis (Sadasivan et al., 2023). In contrast, Liu et al. (2023c) leveraged semantic information, making rule extraction from watermarked texts significantly harder due to varying complexities.
@@
In the context of public detection scenarios, resisting watermark forgeries becomes significantly more challenging. This difficulty arises because attackers can access the watermark detectors. For methods where the watermark generation details are involved in detection (i.e. the hash key in KGW) (Kirchenbauer et al., 2023a, b; Zhao et al., 2023a; Liu et al., 2023c), exposing the watermark generator leads to a complete inability to resist watermark forgeries. To address this issue, Fairoze et al. (2023) have utilized digital signature technology from the field of cryptography. This approach involves generating watermarks using a private key and verifying them with a public key. However, the verification via public key relies on features extracted from the text and users can still exploit these features to forge watermarks. Further advancing this field, Liu et al. (2023b) proposed the use of neural networks for watermark detection. Due to the black-box nature of neural networks, the details of watermark generation are not exposed, which could defend against watermark forgeries in public detection scenarios.